A real estate investment has the following expected cash flows:

                                  Year           Cash Flows

                                    1             $10,000

                                    2              25,000

                                    3              50,000

                                    4              40,000

The discount rate is 7 percent. What is the investment's present value?
Answer #1

Sol:

Cash Flows = Year 1 - $10,000, Year 2 - $25,000, Year 3 - $$50,000, Year 4 - $40,000

Discount rate = 7%

To determine present value of the project:

Computation of Present value of the project
Year Cash Flow PV factor @7% Present value
1 $10,000 0.9346 $9,345.79
2 $25,000 0.8734 $21,835.97
3 $50,000 0.8163 $40,814.89
4 $40,000 0.7629 $30,515.81
Present Value $102,512.46

Therefore present value of the project will be $102,512.46

Working

А B с D 1 Computation of Present value of the project 2 Year Cash Flow PV factor @7% Present value 31 10000 |=1/(1+7%)^1 =B3*

Net present value can also be determined as:

= 10000 / (1 + 7%)^1 + 25000 / (1 + 7%)^2 + 50000 / (1 + 7%)^3 + 40000 / (1 + 7%)^4

= 9345.79 + 21835.97 + 40814.89 + 30515.81 = $102,512.46